# Assessment of Simulation-based Inference Methods for Stochastic Compartmental Models

### v2

@@ 4 Implementation Of The Methods And Their Comparison
In order to render the simulation based approaches computationally feasible, model simulations were conducted using the efficiency of the Julia programming language and the flexible SDE modeling capabilities of SciML and ModelingToolkit.jl [julia_stochdiffeq, ma2021modelingtoolkit]. To approximate the solution of the SDE, we implemented an Euler-Maruyama scheme that constrains trajectories to stay non-negative, effectively clipping potential negative parts and setting them to zero.
@@
For PF, existing software packages were unsuitable as they either lacked Julia support or restricted model specification to simulator-only interfaces. We therefore implemented a custom particle filter in Julia, drawing inspiration from the Python package Particles [chopin_2020]. As an outer sampling scheme, we employed an adaptive Metropolis–Hastings algorithm [adaptive_mh], using the implementation in pyPESTO [pypesto], to account for potential multimodality in the posterior distribution. Missing data were handled by setting the likelihood contribution of the corresponding observation to zero. To mitigate particle degeneracy, we set the number of particles to 200, ensuring that the variance of the likelihood estimate remained below one for parameters around the posterior mode (Supplementary Figure S3). For the outer MH algorithm, we ran 4 Markov chains with a length of 50,000 samples each. We then discarded a burn-in period of the first 25,000 samples per chain. After that, convergence was evaluated using rank‑normalized Gelman–Rubin statistics ($\hat{R}$) and the ESS with autocorrelations up to a maximal lag of 250, both from MCMCDiagnosticTools.jl [turing_jl]. Values of $\hat{R}$ close to 1.0 indicate good convergence and well‑mixed chains [vehtari2021rank]. The ESS assesses and diagnoses poor convergence in the bulk of the distribution due to trends or different locations of the chains. For a given estimand, it is recommended that ESS should exceed $100\cdot\text{nchains}$ [vehtari2021rank]. Together, $\hat{R}$ and ESS provide complementary diagnostics of chain convergence and posterior exploration quality.
@@
For CNF, posterior approximations were obtained via neural posterior estimation with learned summary statistics [radev2020bayesflow], using the BayesFlow software (v1) for amortized Bayesian workflows [radev2023bayesflow]. We trained an independent neural network for each experiment rather than a single globally amortized model. To ensure comparability, all experiments used identical hyperparameters and simulation budgets, determined through a preliminary hyperparameter search. The dimension of the learned summary statistics was set to $2n_{\theta}+2$, where $n_{\theta}$ denotes the number of model parameters. Training was conducted offline with 100,000 simulations for the training dataset and 400 simulations for validation. Models were trained for 100 epochs with a batch size of 32. Summary statistics were extracted using a SequenceNetwork [radev2021outbreakflow] with 64 LSTM units, and the conditional normalizing flow was modeled using an InvertibleNetwork with spline coupling and eight coupling layers, both implemented within the BayesFlow framework. Detailed configurations are provided in Supplementary Information A.5. Handling non‑equidistant time points and missing data required a specialized strategy. We adopted the approach of [wang2024missing], which augments the input with binary indicators marking the presence or absence of each entry. This encoding allows the neural network to account for missing values during inference. For experiments with dense, equidistant data, this additional encoding was not applied. For sparse datasets, the resulting neural networks were trained on the specific missingness pattern present in the simulated data; no amortization across different missingness structures was performed. Empirical cumulative distribution function (ECDF) plots were obtained to assess calibration, and parameter recovery diagnostics were used to evaluate the correspondence between inferred and true parameter values. These diagnostics provide insight into whether the amortized posterior captures uncertainty appropriately and whether point estimates remain consistent with the generative parameters.
@@
For the comparison of both methods we used visual and quantitative measures. The visual comparison using violin and density plots of the marginal posteriors allows comparison of the shape and variability of the posterior approximations. This is complemented by a visual comparison of the model fit. For the evaluations we draw 10,000 parameter vector samples from each approximate posterior distribution and, for each parameter vector, simulate an observable trajectory to create an ensemble. We then plot the central 50%, 90%, and 95% intervals of the ensemble against the data. These are independent posterior predictive samples, obtained by evaluating $p(y\mid\theta)$ for draws $\theta\sim p(\theta\mid\mathcal{D}^{*})$. For a quantitative measure of inference accuracy, we compute the maximum a-posteriori (MAP) estimate in the parameter space via Gaussian kernel density estimation using Silverman’s rule [silverman1998density] as well as the root-mean squared error (RMSE). Specifically, the RMSE was calculated for trajectories simulated from each of the 10,000 posterior samples and then averaged, providing an estimate of the expected RMSE under the posterior distribution. For synthetic experiments, the RMSE was also computed with respect to the true parameters used to generate the data. For the real-data experiment, true parameters are not available. Instead, we report RMSE relative to published parameter estimates from [ethiopian_data], which serve as a reference baseline.
@@ 5 Results
We examined inference with CNF and PF for stochastic epidemic models through a sequence of in silico experiments on synthetic data, following the workflow in Figure 1(c). We begin with the standard SIR model (Section 2.1) using a synthetic prior to validate both methods. We then turn to the two-variant SEIR model (Section 2.2) with literature-informed priors [mcaloon2020incubation, fang2020comparisons], designed to probe the influence of parameter non-identifiability (Section A.3). Within the two-variant SEIR setting, we additionally study the effects of data sparsity and timing mismatch and investigate a reparametrization aimed at reducing non-identifiability. For completeness, Table 3 summarizes all settings and variations. Finally, we apply both approaches to longitudinal data from Ethiopia [ethiopian_data] to assess performance under realistic conditions (Section 5.4). Additional configurations and results are provided in the Supplement (Supplementary Tables S4S18).
@@
in Ethiopia [ethiopian_data].
@@
We compared convergence and posterior estimates obtained with CNF and PF on the SIR model, using two distinct simulated dense datasets (details in Supplementary Information A.6). The prior distributions and parameter values are summarized in Supplementary Tables S3 and S4.
Both CNF and PF yielded highly consistent posterior estimates for the SIR model. The marginal and joint posterior distributions overlapped almost perfectly (Figure 2A), and posterior predictive checks showed that epidemic trajectories generated from either method fit the observed data equally well (Figure 2B). MAP estimates from CNF and PF were nearly identical and close to the true parameters, with RMSEs for infection counts and seroprevalence matching those obtained from simulations with the ground truth (Table 4).
Beyond these direct comparisons, the methods differed in sampling quality. PF achieved excellent convergence, with ESS exceeding 2000 and $\mathbf{\hat{R}}$ values close to 1, indicating robust posterior exploration and well-mixed chains (Supplementary Tables S6 and S7). CNF, in contrast, showed deviations in SBC calibration diagnostics (Supplementary Figure S23), reflecting limitations of the amortized posterior. Notably, these calibration issues were not apparent in the parameter recovery plots: recovered parameters aligned closely with the ground truth and error bars remained narrow (Supplementary Figure S27). This highlights that while CNF’s posterior uncertainty is imperfectly calibrated in a formal sense, the practical impact on point estimates and predictive accuracy is minimal.
@@
In summary, CNF and PF both provide robust Bayesian inference for small stochastic compartmental models. Across datasets, both methods produced equivalent posterior distributions and predictive performance. Notably, CNF, despite being an amortized approach, achieved greater computational efficiency in our experiments, roughly 10-times (Supplementary Table S27), while PF excelled in sampling diagnostics. Moreover, since CNF relies purely on simulations it has consistent computational costs across different parameter vectors. In contrast, PF execution times vary greatly, depending on the posterior geometry. Additionally, the evaluation of the likelihood can be computationally expensive in regions where outcomes are highly sensitive to parameter changes or when parameter vectors do not fit the data well.
Table 4: MAP and RMSE for SIR model.
$\mathbf{\beta}$ $\mathbf{\gamma^{-1}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
True Par. 0.1000 20.00 0.0193 0.0230
CNF 0.1026 18.99 0.0168 0.0201
PF 0.1021 19.32 0.0150 0.0121
@@ 5.2 Non-Identifiability Issues Lead To Posterior Shape Differences
Previous studies have shown that parameters in common epidemiological models are often weakly determined by commonly available datasets [raimundez2021covid]. To assess how such non-identifiabilities affect inference reliability, we compared PF and CNF on the introduced two‑variant SEIR-model with pronounced non-identifiability, making it a suitable test case for evaluating how posterior estimates differ under CNF and PF.
We first examined posterior distributions and predictive performance using four dense simulated observational time series. Posterior distributions showed markedly lower variance compared to the corresponding priors, indicating strong information gain from the data (Figure 3). The notable exception is $\kappa^{-1}$ under CNF, whose marginal remained close to the prior. This behavior is typical in models where certain parameters are weakly informed by the data (see [wang2021posterior] for a more thorough discussion). The PF posterior approximation showed a higher contraction than that of CNF. Additionally, it was also largely enveloped in the CNF posterior approximation, across parameters (Figure 4A). Despite these width differences, epidemic trajectories generated from both methods fit the data equally well (Figure 4B), and MAP estimates together with RMSE scores were close to the true parameters (Table 5). PF estimates were marginally closer to the ground truth, highlighting the method’s focus on trajectory agreement through resampling.
@@
As a secondary analysis, we assessed calibration and convergence diagnostics. CNF showed good calibration in ECDF plots (Supplementary Figure S20), though parameter recovery was imperfect for several parameters and markedly worse for $\kappa^{-1}$, which showed weak correlation with the ground truth (Supplementary Figure S26). PF chains displayed good convergence, with $\mathbf{\hat{R}}$ being close to one across parameters (Supplementary Table S14). However, ESS were relatively low ($\sim$500–600) for several parameters (Supplementary Table S13), indicating limited exploration of posterior tails.
@@
Table 5: MAP and RMSE for full SEIR model.
$\boldsymbol{\gamma^{-1}}$ $\boldsymbol{\kappa^{-1}}$ $\boldsymbol{\beta}$ $\mathbf{s}$ $\mathbf{t_{\mathrm{var}}}$ $\mathbf{I_{0}}$ $\mathbf{RMSE_{infc}}$ $\mathbf{RMSE_{prev}}$
True Par. 17.00 5.00 0.0800 3.00 150.00 500.00 0.0200 0.0245
CNF 15.47 4.93 0.0867 3.29 160.59 446.30 0.0237 0.0237
PF 15.62 5.04 0.0855 3.22 153.02 479.16 0.0228 0.0291
In practice, observations are often non‑equidistant or partially missing. To reflect this, we conducted additional experiments with reduced observation density (Section 2.3). Results for sparse datasets were consistent with those obtained under dense observations. Posterior distributions from CNF and PF showed similar shapes, and epidemic trajectories generated from both methods fit the data equally well (Supplementary Tables S8 to S13), with MAP estimates and RMSE values comparable to the dense setup (Supplementary Table S16). Additionally, CNF remained well calibrated and PF continued to show well‑mixed chains, though ESS values remained low for several parameters, mirroring the diagnostics observed in the dense case (Supplemetary Tables S17 and S18). Together, these results indicate that both methods can accommodate sparse and irregular data without loss of inference quality, though CNF requires careful amortization design to handle heterogeneous observation schedules.
Taken together, these results show that CNF and PF both provide accurate inference for the two‑variant SEIR model, even under pronounced non‑identifiability. CNF yields broader posterior approximations that better capture low‑probability regions, but do not strongly support specific parameter values. PF, in contrast, produces narrower posteriors and more certain point estimates, despite limited exploration of posterior tails due to low ESS. Predictive performance was high for both methods across dense and sparse datasets, with consistent MAP and RMSE values. Sparse observations introduced implementation overhead for CNFs but did not substantially affect inference quality for either method. Additionally, CNF were approximately 10-times faster independent of the used dataset (Supplementary Table S27).

### v3

@@ 4 Implementation Of The Methods And Their Comparison
In order to render the simulation based approaches computationally feasible, model simulations were conducted using the efficiency of the Julia programming language and the flexible SDE modeling capabilities of SciML and ModelingToolkit.jl [Rackauckas and Nie, 2017; Ma et al., 2021]. To approximate the solution of the SDE, we implemented an Euler-Maruyama scheme that constrains trajectories to stay non-negative, effectively clipping potential negative parts and setting them to zero.
@@
For PF, existing software packages were unsuitable as they either lacked Julia support or restricted model specification to simulator-only interfaces. We therefore implemented a custom particle filter in Julia, drawing inspiration from the Python package Particles [Chopin and Papaspiliopoulos, 2020]. As an outer sampling scheme, we employed an adaptive Metropolis–Hastings algorithm [Haario et al., 2001], using the implementation in pyPESTO [Schälte et al., 2023], to account for potential multimodality in the posterior distribution. Missing data were handled by setting the likelihood contribution of the corresponding observation to zero. To mitigate particle degeneracy, we set the number of particles to 200, ensuring that the variance of the likelihood estimate remained below one for parameters around the posterior mode (Supplementary Fig. S1.1). For the outer MH algorithm, we ran 4 Markov chains with a length of 50,000 samples each. We then discarded a burn-in period of the first 25,000 samples per chain. After that, convergence was evaluated using rank‑normalized Gelman–Rubin statistics ($\hat{R}$) and the ESS with autocorrelations up to a maximal lag of 250, both from MCMCDiagnosticTools.jl [Fjelde et al., 2025]. Values of $\hat{R}$ close to 1.0 indicate good convergence and well‑mixed chains [Vehtari et al., 2021]. The ESS assesses and diagnoses poor convergence in the bulk of the distribution due to trends or different locations of the chains. For a given estimand, it is recommended that ESS should exceed $100\cdot\text{nchains}$ [Vehtari et al., 2021]. Together, $\hat{R}$ and ESS provide complementary diagnostics of chain convergence and posterior exploration quality.
@@
For CNF, posterior approximations were obtained via neural posterior estimation with learned summary statistics [Radev et al., 2020], using the BayesFlow software (v1) for amortized Bayesian workflows [Radev et al., 2023]. We trained an independent neural network for each experiment rather than a single globally amortized model. This preserves amortization over latent trajectories and parameter values within each experiment, while avoiding the need to amortize over continuously varying noise levels across experiments, a setting in which global amortization would be inefficient and unnecessary. The resulting per‑experiment amortization still enables efficient SBC diagnostics and fast posterior evaluation once training is complete. Rather than fine‑tuning each method for every individual experiment, which would confound performance differences with tuning effort, we selected CNF hyperparameters through a principled search procedure that we repeated to assess stability. We verified that training was well‑behaved by inspecting training and validation loss curves, which consistently stabilized well before 100 epochs. The dimension of the learned summary statistics was set to $2n_{\theta}+2$, where $n_{\theta}$ denotes the number of model parameters. Training was conducted offline with 100,000 simulations for the training dataset and 400 simulations for validation. Models were trained for 100 epochs with a batch size of 32. Summary statistics were extracted using a SequenceNetwork [Radev et al., 2021] with 64 LSTM units, and the conditional normalizing flow was modeled using an InvertibleNetwork with spline coupling and eight coupling layers, both implemented within the BayesFlow framework. Detailed configurations are provided in Supplementary Information S1.1.3. Handling non‑equidistant time points and missing data required a specialized strategy. We adopted the approach of Wang et al. [2024], which augments the input with binary indicators marking the presence or absence of each entry. This encoding allows the neural network to account for missing values during inference and provides a consistent representation when combining observation functions recorded at different time points. In our setup, only the dimensional consistency of the two observation functions is strictly required; however, we included the binary mask deliberately to facilitate rapid amortization for practitioners who may wish to extend our framework. For the present experiments, this amortization capability was not used: for sparse datasets, the neural networks were trained solely on the specific missingness pattern present in the simulated data to ensure a fair comparison between PF and CNF and to avoid confounding method differences with amortization artifacts. For experiments with dense, equidistant data, this additional encoding was not applied. Empirical cumulative distribution function (ECDF) plots were obtained to assess calibration, and parameter recovery diagnostics were used to evaluate the correspondence between inferred and true parameter values. These diagnostics provide insight into whether the amortized posterior captures uncertainty appropriately and whether point estimates remain consistent with the generative parameters.
@@
For the comparison of both methods we used visual and quantitative measures. The visual comparison using density plots of the marginal posteriors allows comparison of the shape and variability of the posterior approximations. This is complemented by a visual comparison of the model fit. For the evaluations we draw 10,000 parameter vector samples from each approximate posterior distribution and, for each parameter vector, simulate an observable trajectory to create an ensemble. We then plot the central 50%, 90%, and 95% intervals of the ensemble against the data. These are independent posterior predictive samples, obtained by evaluating $p(y\mid\theta)$ for draws $\theta\sim p(\theta\mid\mathcal{D}^{*})$. As an additional qualitative indicator, we report the maximum a‑posteriori (MAP) estimate computed via Gaussian kernel density estimation using Silverman’s rule [Silverman, 1998], which provides a consistent definition across methods. For a quantitative measure of inference accuracy, we first assess marginal coverage by reporting how frequently the true generative parameters fall within the 95% marginal posterior intervals across datasets. We also quantify posterior differences using the 1‑Wasserstein distance in log‑parameter space. All Wasserstein distances are obtained by repeatedly subsampling 1,000 draws per posterior, and we report in‑distribution values to characterize the Monte Carlo error of this procedure. Heuristically, the 1‑Wasserstein distance in log‑parameter space measures the average multiplicative shift required to transform one posterior into another, providing an interpretable notion of scale‑adjusted discrepancy. Concretely, we compute these distances in three settings:
  • Baseline comparisons: when a baseline is available, we measure the Wasserstein distance between each method and the baseline and compare it to the corresponding in‑distribution reference.
  • Model‑parameterization comparison: for the full and reparametrized SEIR2v models, we evaluate how the Wasserstein distance changes when keeping the dataset fixed, again reporting the in‑distribution values as a baseline.
  • Within‑method variability: for one representative parameter set, we assess the variability of each method by computing Wasserstein distances between multiple PF runs or multiple CNF trainings.
Finally, we assess predictive accuracy via the energy score of the posterior predictive distributions. For synthetic experiments, the predictive energy score was also computed with respect to the true parameters used to generate the data. For the real-data experiment, true parameters are not available. Instead, we report the energy score relative to published parameter estimates from [Gudina et al., 2021], which serve as a reference baseline.
@@ 5 Results
We examined inference with CNF and PF for stochastic epidemic models through a sequence of in-silico experiments on synthetic data, following the workflow in Figure 1(d). To assess the performance of both inference methods against a baseline posterior distribution, we use the standard SIS (Section 2.1) and SIR (Section 2.2) model with a synthetic prior. The baseline posterior was computed on a finely discretized version of the SDE using a state-of-the-art Hamiltonian Monte Carlo sampler [Hoffman et al., 2014](Details in Supplementary Information S1.1.4). We then turn to the complex two-variant SEIR model (Section 2.3) with literature-informed priors [McAloon et al., 2020; Fang et al., 2020], designed to probe the influence of parameter non-identifiability. Within the two-variant SEIR setting, we additionally study the effects of data sparsity and timing mismatch and investigate a reparametrization aimed at reducing non-identifiability. For completeness, Table 4 summarizes all settings and variations. Finally, we apply both approaches to longitudinal data from Ethiopia [Gudina et al., 2021] to assess performance under realistic conditions. Evaluation metrics are described in Supplementary Information S1.2 and synthetic dataset generation together with model setups are described in detail in Supplementary Information S1.3.
@@
in Ethiopia [Gudina et al., 2021].
5.1 CNF and PF Agree With A Reference Posterior For The SIS Model
We evaluated convergence and posterior estimates for the SIS model using CNF, PF, and a Hamiltonian Monte Carlo (HMC) baseline applied to a finely discretized version of the SDE. Across 10 synthetic dense datasets, all methods achieved excellent marginal coverage and produced highly consistent Wasserstein distances. Two datasets (sis‑2 and sis‑4) exhibited unusually large distances and were excluded from the aggregate analysis but even for these outliers, visual posterior differences remained minor.
After exclusion, the in‑distribution Monte Carlo baseline (HMC–HMC) was small, with values ranging from 0.056 to 0.119. Between‑method distances were only moderately larger: CNF–PF ranged from 0.063 to 0.156, CNF–HMC ranged from 0.065 to 0.176 and PF–HMC ranged from 0.058 to 0.124. Interpreted on the log‑parameter scale, these values correspond to average multiplicative shifts of roughly $6\%-13\%$ for the baseline and $6\%-17\%$, $7\%-19\%$, and $6\%-13\%$ for the between‑method comparisons, respectively. These results indicate that systematic differences between inference algorithms are small, consistent with the strong visual overlap in marginal and joint posteriors (Figure 2A). Posterior predictive simulations from all three methods reproduced the observed epidemic trajectories with comparable accuracy (Figure 2B), as evidenced by consistent energy scores computed from the posterior predictive distributions of the methods and the ground truth (Table 5).
In terms of sampling behavior, PF showed excellent convergence, with ESS values above 2000 and $\mathbf{\hat{R}}$ near 1. CNF showed good calibration in simulation‑based calibration diagnostics, with near‑uniform rank histograms and accurate parameter recovery. Overall, both PF and CNF matched the predictive and inferential performance of the HMC reference.
Full diagnostics, including ESS, $\mathbf{\hat{R}}$, Wasserstein distance results, calibration plots, and parameter-recovery analyses are provided in Supplementary Information S2.
Figure 2: Evaluation of the posterior approximations for the SIS model
Table 5: Predictive Energy Score for the SIS model.
sis-1 sis-2 sis-3 sis-4 sis-5
True $0.034$ $0.051$ $0.028$ $0.019$ $0.024$
CNF $0.035$ $0.046$ $0.025$ $0.018$ $0.024$
PF $0.034$ $0.046$ $0.025$ $0.018$ $0.023$
HMC $0.035$ $0.046$ $0.025$ $0.018$ $0.023$
sis-6 sis-7 sis-8 sis-9 sis-10
True $0.041$ $0.025$ $0.045$ $0.042$ $0.041$
CNF $0.040$ $0.027$ $0.044$ $0.035$ $0.040$
PF $0.037$ $0.022$ $0.042$ $0.031$ $0.040$
HMC $0.037$ $0.022$ $0.042$ $0.031$ $0.039$
@@ 5.1 CNF and PF Agree With A Reference Posterior For The SIS Model
We compared convergence and posterior estimates obtained with CNF and PF on the SIR model, using twelve distinct synthetic dense datasets. As a reference, we additionally computed a reference posterior using Hamiltonian Monte Carlo (HMC) applied to a finely discretized version of the underlying SDE.
The in‑distribution Monte Carlo baseline (HMC–HMC) ranged from 0.034 to 0.096. Between‑method distances were modestly larger: CNF–PF ranged from 0.063 to 0.165, CNF–HMC from 0.076 to 0.183, and PF–HMC from 0.048 to 0.125. Interpreted on the log‑parameter scale, these values correspond to average multiplicative shifts of approximately 3–10% for the baseline and 6–18%, 8–20%, and 5–13% for the between‑method comparisons, respectively. Posterior predictive checks showed that epidemic trajectories generated from all three methods fit the observed data equally well (Figure 3B), as evidenced by the close agreement in predictive energy scores (Table 6).
Across all datasets, PF and CNF achieved 83.33% marginal coverage and HMC achieved 87.5%, and the three posterior approximations showed high agreement overall. Marginal and joint posterior distributions overlapped closely (Figure 3A) with CNF exhibiting a small systematic offset to HMC and PF.This offset is consistent with the mild posterior miscalibration discussed below. The Wasserstein distances reflect this pattern. The in‑distribution Monte Carlo baseline (HMC–HMC) ranged from 0.034 to 0.096. Between‑method distances were modestly larger: CNF–PF ranged from 0.063 to 0.165, CNF–HMC from 0.076 to 0.183 and PF–HMC ranged from 0.048 to 0.125. Interpreted on the log‑parameter scale, these values correspond to average multiplicative shifts of approximately 3–10% for the baseline and 6–18%, 8–20%, and 5–13% for the between‑method comparisons, respectively. Detailed results are provided in Supplementary Fig. S3.13. Posterior predictive checks showed that epidemic trajectories generated from all three methods fit the observed data equally well (Figure 3B), as evidenced by the close agreement in predictive energy scores (Table 6).
Beyond these direct comparisons, the methods differed in sampling quality. PF achieved excellent convergence, with ESS exceeding 2000 and $\mathbf{\hat{R}}$ values close to 1, indicating robust posterior exploration and well-mixed chains (Supplementary Tables S3.2 and S3.3). CNF, in contrast, showed deviations in SBC diagnostics (Supplementary Fig. S3.14), reflecting limitations of the amortized posterior. Notably, these calibration issues were not apparent in the parameter recovery plots: recovered parameters aligned closely with the ground truth and error bars remained narrow (Supplementary Fig. S3.16). The miscalibration geometry also matches the observed offset of CNF relative to PF and HMC; for example, an upward‑sloping SBC histogram for $\beta$ corresponds to a left-shift in the $\beta$ marginal. This highlights that while CNF’s posterior uncertainty is imperfectly calibrated in a formal sense, the practical impact on point estimates and predictive accuracy is minimal.
@@
In summary, CNF and PF both provide robust Bayesian inference for small stochastic compartmental models. Across datasets, both methods produced equivalent posterior distributions and predictive performance. Notably, CNF, despite being an amortized approach, achieved greater computational efficiency in our experiments (roughly 10-fold) while PF excelled in sampling diagnostics. Moreover, since CNF relies purely on simulations it has consistent computational costs across different parameter vectors. In contrast, PF execution times vary greatly, depending on the posterior geometry. Additionally, the evaluation of the likelihood can be computationally expensive in regions where outcomes are highly sensitive to parameter changes or when parameter vectors do not fit the data well.
Full diagnostics, including ESS, $\mathbf{\hat{R}}$, Wasserstein distance results, calibration plots, and parameter-recovery analyses are provided in Supplementary Information S3.
Table 6: Predictive Energy Score for the SIR model.
sir-1 sir-2 sir-3 sir-4 sir-5 sir-6
True $0.046$ $0.057$ $0.032$ $0.020$ $0.052$ $0.021$
CNF $0.046$ $0.040$ $0.033$ $0.033$ $0.052$ $0.018$
PF $0.045$ $0.039$ $0.033$ $0.025$ $0.054$ $0.020$
HMC $0.045$ $0.039$ $0.032$ $0.024$ $0.054$ $0.021$
sir-7 sir-8 sir-9 sir-10 sir-11 sir-12
True $0.050$ $0.054$ $0.056$ $0.071$ $0.061$ $0.073$
CNF $0.048$ $0.053$ $0.057$ $0.069$ $0.061$ $0.066$
PF $0.047$ $0.052$ $0.057$ $0.070$ $0.062$ $0.060$
HMC $0.047$ $0.053$ $0.057$ $0.069$ $0.061$ $0.060$
@@ 5.3 Non-Identifiability Issues Lead To Posterior Shape Differences
Previous studies have shown that parameters in common epidemiological models are often weakly determined by commonly available datasets Raimúndez et al. [2021]. To assess how such non-identifiabilities affect inference reliability, we compared PF and CNF on the introduced two‑variant SEIR-model with pronounced non-identifiability for specific parameter combinations.
We first examined posterior distributions and predictive performance using twelve dense simulated observational time series. Posterior distributions showed markedly lower variance compared to the corresponding priors, indicating strong information gain from the data (Figure 4). The notable exception is $\kappa^{-1}$ under CNF, whose marginal remained close to the prior for some experiments, consistent with weak identifiability Wang et al. [2021]. PF posteriors were generally more concentrated than those of CNF and were largely enveloped by the CNF approximation across parameters (Figure 5A). Despite these width differences, both methods generally include the true parameters in their posteriors, with CNF achieving 95.2% marginal coverage and PF achieving 91.7%. Epidemic trajectories generated from both methods fit the data equally well (Figure 5B) with comparable energy scores (Table 7). PF produced marginally lower energy scores, highlighting the method’s focus on trajectory agreement through resampling.
@@
As an additional diagnostic, Wasserstein distances showed that variability across repeated CNF runs (CNF-CNF) was of the same order as the differences between CNF and PF (CNF-PF) indicating that the two methods agree up to the intrinsic stochasticity of the CNF estimator. CNF showed good calibration in ECDF plots, though parameter recovery was imperfect for several parameters and markedly worse for $\kappa^{-1}$, which showed weak correlation with the ground truth. PF chains displayed good convergence, with $\mathbf{\hat{R}}$ being close to one across parameters. However, ESS were relatively low (<1000<1000) for several parameter sets, indicating limited exploration of posterior tails.
@@
Table 7: Predictive Energy Score for the full SEIR2V model with dense data.
d-1-1 d-1-2 d-2-1 d-2-2 d-3 d-4 d-5
True $0.113$ $0.115$ $0.090$ $0.112$ $0.059$ $0.126$ $0.089$
CNF $0.108$ $0.106$ $0.090$ $0.098$ $0.058$ $0.124$ $0.060$
PF $0.107$ $0.104$ $0.088$ $0.098$ $0.057$ $0.123$ $0.059$
d-6 d-7 d-8 d-9 d-10 d-11 d-12
True $0.118$ $0.100$ $0.101$ $0.108$ $0.105$ $0.099$ $0.120$
CNF $0.116$ $0.109$ $0.087$ $0.102$ $0.103$ $0.088$ $0.118$
PF $0.115$ $0.114$ $0.088$ $0.101$ $0.097$ $0.086$ $0.117$
In practice, observations are often non‑equidistant or partially missing. To reflect this, we conducted additional experiments with reduced observation density (Section 2.4). Results for sparse datasets were consistent with those obtained under dense observations. Posterior distributions from CNF and PF showed similar shapes, and epidemic trajectories generated from both methods fit the data equally well, with energy scores of comparable magnitude (Table 8). Additionally, CNF remained well calibrated and PF continued to show well‑mixed chains, though ESS values remained low for several parameters, mirroring the diagnostics observed in the dense case. Together, these results indicate that both methods can accommodate sparse and irregular data without loss of inference quality, though CNF requires careful amortization design to handle heterogeneous observation schedules.
Table 8: Predictive Energy Score for the full SEIR2V model with sparse data.
s-1-1-1 s-1-1-2 s-1-1-3 s-1-2-1 s-1-2-2 s-1-2-3 s-3 s-4
True $0.042$ $0.039$ $0.039$ $0.041$ $0.053$ $0.038$ $0.018$ $0.056$
CNF $0.039$ $0.040$ $0.035$ $0.041$ $0.036$ $0.033$ $0.015$ $0.038$
PF $0.037$ $0.038$ $0.032$ $0.033$ $0.035$ $0.033$ $0.015$ $0.038$
s-5 s-6 s-7 s-8 s-9 s-10 s-11 s-12
True $0.030$ $0.037$ $0.029$ $0.035$ $0.052$ $0.036$ $0.036$ $0.042$
CNF $0.017$ $0.032$ $0.030$ $0.029$ $0.041$ $0.029$ $0.033$ $0.037$
PF $0.015$ $0.032$ $0.030$ $0.029$ $0.038$ $0.024$ $0.030$ $0.036$
Taken together, these results show that CNF and PF both provide accurate inference for the two‑variant SEIR model, even under pronounced non‑identifiability. CNF yields broader posterior approximations that better capture low‑probability regions, but do not strongly support specific parameter values. PF, in contrast, produces narrower posteriors and more certain point estimates, despite limited exploration of posterior tails due to low ESS. Predictive performance was high for both methods across dense and sparse datasets, with consistent MAP and RMSE values predictive energy scores. Sparse observations introduced implementation overhead for CNFs but did not substantially affect inference quality for either method. Additionally, CNF were approximately 10-times faster independent of the used dataset (Supplementary Table S1.22).
Full diagnostics, including calibration plots, convergence statistics, and parameter‑recovery analyses, are provided in Supplementary Information S4 (dense datasets) and S5 (sparse datasets).